OpenStudy (anonymous):

Find the nth term of 25,32,39 and 46 T(n)=a+(n-1)d T(n)=a+(n-1)7 T(n)=25+(n-1)7 What is the next step?

OpenStudy (zarkon):

there isn't a next step

OpenStudy (anonymous):

??

OpenStudy (zarkon):

the nth term is \[a_n=25+(n-1)\cdot 7\]

OpenStudy (zarkon):

n=1,2,...

OpenStudy (anonymous):

what is the (n-1) "n"

OpenStudy (anonymous):

=,='''

OpenStudy (zarkon):

well you lost me there with your notation...are you unhappy that is have an n-1 and not n?

OpenStudy (anonymous):

i dont understand what n stands for i mean a stands for the first term d stands for th common difference what does n stand for?

OpenStudy (zarkon):

n is the term number so if n=11 then \[a_{11}\] is the 11th term in the sequence \[a_{11}=25+(11-1)\cdot 7=25+10\cdot 7=25+70=95\]

OpenStudy (anonymous):

how can u find the n?

OpenStudy (zarkon):

n can be any positive integer

OpenStudy (zarkon):

there really isn't anything to find unless they were to ask something like this for what value of n is \[a_n=95\] the answer would be n=11

OpenStudy (anonymous):

T(n)=25+(n-1)7 so can the next step be T(n)=25+(7n-7) T(n)=25-7+7n

OpenStudy (zarkon):

sure...if you want

OpenStudy (zarkon):

25+(n-1)7 is a perfectly fine nth term though

OpenStudy (anonymous):

so the next would be T(n)=18+7n T(n)=7n+18

OpenStudy (anonymous):

ok...but i need the most simplified one

OpenStudy (zarkon):

then that would be it

OpenStudy (zarkon):

7n+18

OpenStudy (anonymous):

thanks, and by the way, i have became a fan of yours to contact you more easily

OpenStudy (zarkon):

np

OpenStudy (anonymous):

??

OpenStudy (anonymous):

what is np?

OpenStudy (zarkon):

no problem :)

OpenStudy (anonymous):

ok, when is the time you will be here

OpenStudy (zarkon):

I don't know...I don't keep office hours ;)

OpenStudy (anonymous):

oic:)

OpenStudy (anonymous):

do u have a facebook of msn accont?

OpenStudy (zarkon):

this is really late for me...usually I'm not on at this hour

OpenStudy (zarkon):

no

OpenStudy (anonymous):

ok thanks anyway

OpenStudy (anonymous):

:)

OpenStudy (zarkon):

np ;)

OpenStudy (anonymous):

88;)